OpenStudy (anonymous):

Eliminate the parameter. x=t^2+2, y=t^2-4 a. y = x - 6, x ≥ 1 b. y = x + 6, x ≥ 1 c. y = x2 - 6, x ≥ 1 d. y = x2 + 6, x ≥ 1

hartnn (hartnn):

we need to eliminate t^2, right ? so can you isolate t^2 from \( x= t^2+2 \)??

hartnn (hartnn):

hint : subtract 2 from both sides :)

OpenStudy (anonymous):

x-2=t^2

hartnn (hartnn):

yes! now plug that t^2 in y= t^2 -4

hartnn (hartnn):

y = (x-2) -4 simplify :)

OpenStudy (anonymous):

y=-4x+8

hartnn (hartnn):

not actually y = x-2 -4 is just y= x-6

hartnn (hartnn):

option A :)

OpenStudy (anonymous):

Oh yeah!

hartnn (hartnn):

:)

OpenStudy (anonymous):

Can you help me with one more please?

hartnn (hartnn):

for sure :)

OpenStudy (anonymous):

Eliminate the parameter. x = 3 cos t, y = 3 sin t

hartnn (hartnn):

square and add 1) x= 3 cos t Square both sides 2) y= 3 sin t square both sides what do u get ?

OpenStudy (anonymous):

x^2=3 cos^2 t y^2=3 sin^2 t

hartnn (hartnn):

now add these equations!

OpenStudy (anonymous):

How?

hartnn (hartnn):

\(\Large x^2+y^2 =3 cos^2 t +3 sin^2 t \) can you simplify this ?

OpenStudy (anonymous):

I have no idea what to do

hartnn (hartnn):

factor out the 3 from the right!

hartnn (hartnn):

and then use \(\large \sin^2x +\cos^2 x =1 \)

OpenStudy (anonymous):

So where would I have the 3?

hartnn (hartnn):

\(x^2 + y^2 = 3 (\sin^2x +\cos^2 x ) = 3 \times 1 \\ \Large x^2+y^2 = 3 \) thats it!

OpenStudy (anonymous):

Oh! Thank you so much!! :)

hartnn (hartnn):

most welcome ^_^